# Unified Multimodal Understanding and Generation Models: Advances, Challenges, and Opportunities

### v2

@@ 3.3 Fused Autoregressive and Diffusion Models
2) Hybrid Encoding: As shown in Fig. 5 (c-2), hybrid encoding fuses both semantic features (e.g., from CLIP or ViT encoders) and pixel-level latents (e.g., from SD-VAE), providing a more expressive image representation. This approach allows models to leverage high-level semantic abstraction while maintaining detailed visual information. Representative example is Janus-Flow [179] that adopts a dual-encoder architecture and presents a minimalist architecture that harmonizes AR language models with rectified flow. It decouples the understanding and generation encoders, using SigLIP as the vision encoder for multimodal understanding and SDXL-VAE for image generation. However, the pseudo hybrid encoding design limits the model’s ability to simultaneously leverage both semantic and pixel-level features during generation, as only the pixel encoder is active in the image synthesis process. This decoupling, while beneficial for modularity and training efficiency, prevents the model from fully exploiting semantic cues during image decoding, potentially weakening fine-grained alignment and multimodal compositionality in generative tasks.
@@ 4.2 Text-to-Image Datasets
  • Mario-10M [208] and AnyWord-3M [210]: These two datasets focus on accurate text rendering within images. Mario-10M (10M samples), used for training the TextDiffuser model [208], and AnyWord-3M (3M samples), used to train AnyText [210], provide data specifically designed to improve the legibility and placement of text generated in images.
@@ 5 Benchmarks
Interleaved Generation
@@ 5.2 Evaluation on Image Generation
Text‐to‐Image Generation. Early automated metrics such as FID [284] and CLIPScore [22] laid the foundation for image quality assessment. However, recent benchmarks shift the focus toward compositionality, alignment, and real-world applicability. GenEval [263] evaluates six fine-grained tasks, including single-object generation, object co-occurrence, counting, color control, relative positioning, and attribute binding by comparing outputs from pretrained detectors against ground-truth annotations.
@@
To support large-scale concept diversity, EvalMuse-40K [286] provides 40K crowdsourced prompts focusing on nuanced concept representation, and HEIM [265] identifies 12 aspects, including text-image alignment, image quality, aesthetics, originality, reasoning, knowledge, bias, toxicity, fairness, robustness, multilinguality and efficiency. Considering practical needs, FlashEval [287] shrinks the large-scale evaluation set into diverse smaller ones through iterative search to accelerate the benchmark testing. MEMO-Bench [288] introduces a comprehensive benchmark for evaluating the emotional understanding and expression capabilities of T2I models and MLLMs.
@@
Image Editing. Benchmarks for instruction‐guided image editing have grown in scale and scope. MagicBrush [217] is the first large-scale, manually annotated dataset for instruction-guided real image editing that covers diverse scenarios: single-turn, multi-turn, mask-provided, and mask-free editing. HQ-Edit [218] contains approximately 200K high‐resolution edits with computed alignment and coherence scores, allowing quantitatively assessing the quality of image edit pairs using GPT-4V.
Building up on this, I2EBench [272] consolidates 2K+ images and over 4K multi‐step instructions across 16 editing dimensions. EditVAl [270] provides a standardized benchmark with a curated dataset of images annotated for diverse fine-grained edit types and an automated evaluation pipeline using pretrained vision–language models whose scores strongly correlate with human judgments. Emu-Edit [271] encompasses seven instruction-based editing tasks among background alteration, comprehensive changes, style modification, object removal, object addition, localized edits, and texture alterations, providing human-instruction/image pairs along with input/output descriptions. HumanEdit [273] offers 5,751 high-resolution images paired with open-form language instructions across six edit categories: action, add, counting, relation, remove, and replace, with masks and multi-stage human feedback to rigorously benchmark instruction-guided image editing models.
More recently, GEdit-Bench [274] is proposed, which is a real-world image editing benchmark comprising 606 reference image–instruction pairs curated from over 1K user editing examples, designed to comprehensively evaluate practical image editing models.
Other Types of Image Generation. Beyond text-to-image generation and image editing, additional benchmarks probe large-scale conditional and personalized synthesis. MultiGen-20M [226] comprises over 20 million image–prompt–condition triplets sourced from LAION-Aesthetics-V2 [289], enabling comprehensive automated evaluation of alignment under varied visual conditions and the evaluation set with 100-300 imagecondition-prompt triplets for each task.
DreamBench [280] introduces a personalized generation test spanning 30 reference objects paired with curated prompts and human-annotated fidelity judgments. DreamBench++ [281] extends this framework to 150 diverse reference images and 1,350 prompts, employing advanced multimodal language models for automated, human-aligned scoring across concept preservation, compositional fidelity, and stylistic consistency. Together, these datasets offer a coherent spectrum from massive automated benchmarks to focused, human-centric assessments of conditional and subject-driven image generation.
@@ 5.3 Evaluation on Interleaved Generation
Interleaved evaluation benchmarks challenge models to seamlessly alternate between text and image modalities across multiple turns, reflecting realistic dialogue and storytelling scenarios. InterleavedBench [275] is the first benchmark carefully curated for the evaluation of interleaved textand-image generation, featuring a rich array of tasks to cover diverse real-world use cases and evaluating models on text quality, perceptual fidelity, multimodal coherence and helpfulness. Building on this, ISG [277] introduces scene‐graph annotations and a four‐tier evaluation (holistic, structural, block‐level and image‐specific) over 1K samples in eight scenarios and 21 subtasks, enabling fine‐grained assessment of interleaved text–image outputs.
Other benchmarks emphasize open‐domain instruction and end‐to‐end interleaving. OpenING [279] assembles 5K human‐annotated instances across 56 real‐world tasks (e.g. travel guides, design ideation) with IntJudge to test open-ended multimodal generation methods on arbitrary instruction‐driven interleaved generation. In contrast, OpenLEAF [276] gathers 30 open‐domain queries with each written and reviewed by annotators to probe foundational interleaved text–image generation, measuring entity and style consistency via LMM evaluators plus human validation. Finally, MMIE [278] proposes a unified interleaved suite by sampling from 12 fields and 102 subfields, offering a mix of multiple-choice and open-ended question formats to evaluate models in a diverse manner.
@@ 6 Challenges and Opportunities on Unified Models
Currently, at its rudimentary stage, unified multimodal models face several significant challenges that should be addressed to achieve robust and scalable understanding and generation capabilities. First, the high dimensionality of visual and textual data leads to extremely long token sequences. Efficient tokenization and compression strategies are essential to reduce memory and computation costs while preserving representational fidelity. Second, cross-modal attention becomes a performance bottleneck as image resolution and context length increase. Scalable alternatives such as sparse or hierarchical attention mechanisms may potentially mitigate this issue. Third, pretraining datasets often include noisy or biased image–text pairs, particularly for complex image compositions and interleaved image-text data. Reliable data filtering, debiasing, and synthesizing are crucial to ensure fairness and robustness. Fourth, evaluation protocols are typically designed for single tasks in isolation. There is a growing need for comprehensive benchmarks that assess both understanding and generation in an integrated manner, especially for sophisticated tasks such as image editing and interleaved image-text generation.
To the best of our knowledge, most of current unified multimodal models primarily emphasize image understanding and text-to-image generation, while capabilities such as image editing are only attained through post-finetuning. Moreover, advanced functionalities like spatially controlled image generation, subject(s)-driven image generation, and interleaved image-text generation remain largely unexplored in the unified framework. Consequently, we believe there are abundant opportunities to advance the field by addressing key areas such as architectural design, training efficiency, dataset curation, and evaluation methodologies to achieve unified multimodal models.

### v3

@@ 3.3 Fused Autoregressive and Diffusion Models
2) Hybrid Encoding: As shown in Fig. 5 (c-2), hybrid encoding fuses both semantic features (e.g., from CLIP or ViT encoders) and pixel-level latents (e.g., from SD-VAE), providing a more expressive image representation. This approach allows models to leverage high-level semantic abstraction while maintaining detailed visual information. Specifically, Janus-flow [189], Mogao [190] and BAGEL [191] adopt a dual-encoder architecture and presents a minimalist architecture that harmonizes AR language models with rectified flow. They decouples the understanding and generation encoders, using SigLIP or the concatenation of SigLIP and SDXL-VAE as the vision encoder for multimodal understanding and SDXL-VAE or FLUX-VAE for image generation. However, the pseudo hybrid encoding design limits the model’s ability to simultaneously leverage both semantic and pixel-level features during generation, as only the pixel encoder is active in the image synthesis process. This decoupling, while beneficial for modularity and training efficiency, prevents the model from fully exploiting semantic cues during image decoding, potentially weakening fine-grained alignment and multimodal compositionality in generative tasks.
@@
TokLIP [151] b-1 Qwen2.5 VQGAN+SigLIP VQGAN Causal 2025-05
Selftok [153] b-1 LLaMA3.1 SD3-VAE+MMDiT SD3 Causal 2025-05
@@
DualToken [165] b-2 Qwen2.5 SigLIP RQVAE Causal 2025-03
BLIP3-o [159] b-2 Qwen2.5-VL OpenAI-CLIP Lumina-Next Causal 2025-05
@@
Mogao [190] c-2 Qwen2.5 SigLIP+SDXL-VAE SDXL-VAE SDXL-VAE Bidirect. 2025-05
BAGEL [191] c-2 Qwen2.5 SigLIP FLUX-VAE FLUX-VAE Bidirect. 2025-05
@@ 4.1 Multimodal Understanding Datasets
  • LLaVA-OneVision [216]: This visual instruction tuning collection features two main parts: a Single-Image dataset of 3.2 million diverse, categorized samples (QA, OCR, math, etc.), and the OneVision dataset with 1.6 million mixed-modal samples (including video, multi-image, and selected single-image data).
  • Infinity-MM [216]: Infinity-MM is a comprehensive multimodal training dataset with over 40 million samples, created by extensively collecting and categorizing existing open-source datasets alongside newly generated data. This collection includes image captions, general visual instructions, higher-quality selective instructions, and a significant portion of data generated by GPT-4 or synthesized using a custom VLM-based pipeline to ensure alignment and diversity. All data undergoes rigorous processing and filtering for quality and consistency.
@@ 4.2 Text-to-Image Datasets
  • Text Rendering Datasets: Several datasets have been developed to specifically address the challenges of accurately and legibly rendering text within generated images. Mario-10M [220], with 10 million samples, was used to train the TextDiffuser model [220], providing data designed to improve text placement and legibility. The RenderedText dataset [221] offers 12 million high-resolution synthetic images of handwritten text, generated with diverse visual attributes, serving as a rich resource for handwritten text understanding and generation. AnyWord-3M [223], containing 3 million samples, is crucial for training models like AnyText [223] and also focuses on enhancing the quality of generated text. Lastly, TextAtlas5M [232] targets dense text generation, incorporating a diverse mix of interleaved documents, synthetic data, and real-world images with longer captions and human annotations to tackle complex text-rich image scenarios.
@@
  • Synthesized Datasets: Specialized datasets for text-to-image synthesis are increasingly created using existing generative models. The text-to-image-2M dataset [228] provides  2 million enhanced text-image pairs for fine-tuning, curated using advanced T2I and captioning models. SFHQ-T2I [230] offers 122K diverse, high-resolution synthetic face images generated by multiple T2I models, ensuring variance and privacy. For entity control, the EliGen TrainSet [231] uses images from a baseline model (FLUX.1-dev) and MLLM-generated prompts for stylistic consistency and detailed annotation. Lastly, BLIP-3o 60k [159] provides 60,000 instruction tuning samples distilled from GPT-4o, covering various categories for diverse training.
@@ 4.5 Other Text+Image-to-Image Datasets
  • X2I-subject-driven [84]: The X2I-subject-driven dataset facilitates subject-driven image generation through two components. The GRIT-Entity dataset is derived from the GRIT [212] dataset by automatically detecting and segmenting objects from images, followed by an optional MS-Diffusion [252] repainting step to improve quality and diversity. To encourage more robust generative capabilities beyond simple copy-paste patterns, the higher-quality Web-Images dataset was constructed by identifying notable individuals through automated text analysis and large language model filtering, scraping their web images, performing automated visual verification to ensure subject accuracy, and then captioning the selected images.
@@
General-Bench [279] $>$700 tasks Ultra Large-scale Eval Varied by Task Arxiv2025
@@
DrawBench [72] 200 Prompts Comprehensive Eval Text Prompt $\rightarrow$ Image NeurIPS2022
PartiPrompts [280] 1600 Prompts Comprehensive Eval Text Prompt $\rightarrow$ Image TMLR2022
PaintSkills [281] $\sim$7K Scenes Compositional Eval Text Prompt $\rightarrow$ Image ICCV2023
@@
TIFA [283] 4081 Prompts QA-based Eval Text Prompt $\rightarrow$ Image ICCV2023
@@
DSG-1k [288] 1060 Prompts Compositional Eval Text Prompt $\rightarrow$ Image ICLR2024
@@
ConceptMix [290] 2100 Prompts Compositional Eval Text Prompt $\rightarrow$ Image NeurIPS2024
@@
EditBench [293] 240 Edits Mask-guided Editing Img. + Ins. + [Mask] $\rightarrow$ Image CVPR2023
@@
Reason-Edit [296] 219 Edits Complex Instruction Editing Image + Instruction $\rightarrow$ Image CVPR2024
@@
IE-Bench [299] 301 Edits Human-aligned Perceptual Eval Image + Instruction $\rightarrow$ Image Arxiv2025
@@
CompBench [301] 3K Edits Complex Instruction Editing Image + Instruction $\rightarrow$ Image Arxiv2025
GIE-Bench [302] 1080 Edits Content-preserving Eval Image + Instruction $\rightarrow$ Image Arxiv2025
Interleaved / Compositional Generation
@@
UniBench [308] 81 fine-grained tags Unified Compositional Eval Prompt → Images + Answer Arxiv2025
@@
VTBench [311] Varied Visual Tokenizer Eval Image $\rightarrow$ Reconstructed Image Arxiv2025
5 Benchmarks
@@ 5.1 Evaluation on Understanding
Moreover, General-Bench [279], an ultra-large benchmark comprising over 700 tasks and 325,800 instances across varied modalities and capabilities, provides a synergy-driven evaluation suite for multimodal generalist models.
@@ 5.2 Evaluation on Image Generation
Text‐to‐Image Generation. Early automated metrics such as FID [314] and CLIPScore [22] established the foundation for evaluating image quality. More recent benchmarks, however, emphasize compositional reasoning, prompt alignment, and real-world applicability. PaintSkills [281], DrawBench [72], and PartiPrompts [280] evaluate core compositional capabilities. GenEval [284] evaluates six fine-grained tasks, including single-object generation, object co-occurrence, counting, color control, relative positioning, and attribute binding by comparing outputs from pretrained detectors against ground-truth annotations.
@@
To support large-scale concept diversity, EvalMuse-40K [316] provides 40K crowdsourced prompts focusing on nuanced concept representation, and HEIM [286] identifies 12 aspects, including text-image alignment, image quality, aesthetics, originality, reasoning, knowledge, bias, toxicity, fairness, robustness, multilinguality and efficiency. Considering practical needs, FlashEval [317] shrinks the large-scale evaluation set into diverse smaller ones through iterative search to accelerate the benchmark testing. MEMO-Bench [318] introduces a comprehensive benchmark for evaluating the emotional understanding and expression capabilities of T2I models and MLLMs. ConceptMix [290] evaluates text-to-image models’ compositional generation ability by sampling k-tuples of visual concepts to construct prompts and automatically verifying concept presence in the resulting images using a strong visual language model. TIFA [283] offers a fine-grained benchmark for evaluating text-to-image faithfulness via visual question answering generated from prompts. To enrich dependencies in question generation for VQA-based evaluation of image–prompt alignment, DSG-1k [288] refines its questions using a multi-level semantic graph.
@@
Image Editing. Benchmarks for instruction‐guided image editing have grown in scale and scope. MagicBrush [234] is a large-scale, manually annotated dataset for instruction-guided real image editing that covers diverse scenarios: single-turn, multi-turn, mask-provided, and mask-free editing. HQ-Edit [235] contains approximately 200K high‐resolution edits with computed alignment and coherence scores, allowing quantitatively assessing the quality of image edit pairs using GPT-4V.
Building on this, I2EBench [297] consolidates over 2K images and 4K multi-step instructions across 16 editing dimensions. EditVal [294] offers a standardized benchmark with fine-grained edit annotations and an automated evaluation pipeline aligned with human judgment. Emu-Edit [295] covers seven editing tasks including background changes, object-level edits, and style modifications, with paired instructions and I/O descriptions. Reason-Edit [296] is a diagnostic benchmark targeting causal and counterfactual reasoning, emphasizing object relations, attribute dependencies, and multi-step inference.
Offering masked input–reference pairs across varied objects, attributes, and scenes, EditBench [293] delivers a diagnostic benchmark for text-guided image inpainting that enables precise evaluation of editing quality. HumanEdit [298] includes 5,751 high-resolution images and open-form instructions spanning six edit types, with annotated masks and multi-stage human feedback. IE-Bench [299] provides a human-aligned benchmark for evaluating text-driven image editing quality with diverse edits and perceptual scores.
More recent benchmarks include GEdit-Bench [300] which features 606 real-world instruction–image pairs, CompBench [301] that decomposes edits into location, appearance, dynamics and object dimensions via large-scale MLLM–and–human collaboration, and GIE-Bench [302] which uses multiple-choice VQA and object-aware masking on over 1,000 examples to evaluate editing accuracy and content preservation.
Other Types of Image Generation. Beyond text-to-image generation and editing, additional benchmarks target conditional and personalized synthesis. MultiGen-20M [243] provides over 20 million image–prompt–condition triplets from LAION-Aesthetics-V2 [319], supporting automated evaluation across diverse visual conditions. DreamBench [309] benchmarks personalized generation using 30 reference objects with curated prompts and human fidelity ratings. DreamBench++ [310] scales this to 150 subjects and 1,350 prompts, using advanced vision–language models for human-aligned scoring of concept preservation, composition, and style. Together, these datasets span large-scale automated and fine-grained human-centric evaluation of conditional generation.
VTBench [311] provides a systematic benchmark for evaluating visual tokenizers in autoregressive image generation across image reconstruction, detail preservation, and text preservation.
@@ 5.3 Evaluation on Interleaved Generation
Interleaved evaluation benchmarks challenge models to seamlessly alternate between text and image modalities across multiple turns, reflecting realistic dialogue and storytelling scenarios. InterleavedBench [303] is a representative benchmark carefully curated for the evaluation of interleaved textand-image generation, featuring a rich array of tasks to cover diverse real-world use cases and evaluating models on text quality, perceptual fidelity, multimodal coherence and helpfulness. Building on this, ISG [305] introduces scene‐graph annotations and a four‐tier evaluation (holistic, structural, block‐level and image‐specific) over 1K samples in eight scenarios and 21 subtasks, enabling fine‐grained assessment of interleaved text–image outputs.
Other benchmarks emphasize open‐domain instruction and end‐to‐end interleaving. OpenING [307] assembles 5K human‐annotated instances across 56 real‐world tasks (e.g. travel guides, design ideation) with IntJudge to test open-ended multimodal generation methods on arbitrary instruction‐driven interleaved generation. In contrast, OpenLEAF [304] gathers 30 open‐domain queries with each written and reviewed by annotators to probe foundational interleaved text–image generation, measuring entity and style consistency via LMM evaluators plus human validation. Finally, MMIE [306] proposes a unified interleaved suite by sampling from 12 fields and 102 subfields, offering a mix of multiple-choice and open-ended question formats to evaluate models in a diverse manner. In a more recent work, UniBench [308] was introduced as a comprehensive compositional benchmark for evaluating unified models, offering 81 fine-grained tags to ensure high diversity.
@@ 6 Challenges and Opportunities on Unified Models
Currently, at its rudimentary stage, unified multimodal models face several significant challenges that should be addressed to achieve robust and scalable understanding and generation capabilities. First, the high dimensionality of visual and textual data leads to extremely long token sequences. Efficient tokenization and compression strategies are essential to reduce memory and computation costs while preserving representational fidelity. Second, cross-modal attention becomes a performance bottleneck as image resolution and context length increase. Scalable alternatives such as sparse or hierarchical attention mechanisms may potentially mitigate this issue. Third, pretraining datasets often include noisy or biased image–text pairs, particularly for complex image compositions and interleaved image-text data. Reliable data filtering, debiasing, and synthesizing are crucial to ensure fairness and robustness. Fourth, evaluation protocols are typically designed for single tasks in isolation. There is a growing need for comprehensive benchmarks that assess both understanding and generation in an integrated manner, especially for sophisticated tasks such as image editing and interleaved image-text generation. Apart from the issues and challenges on architecture, data, and evaluation, applying chain-of-thought (CoT) reasoning and reinforcement learning (RL) techniques into unified MLLM models to improve both interpretability and performance [320] is also worth exploring. CoT can guide the model to generate intermediate reasoning steps, which are particularly beneficial for complex visual-question answering or image-conditioned generation. Meanwhile, RL can be used to optimize long-horizon objectives such as factual consistency, user satisfaction, or task success rate beyond token-level likelihoods. Moreover, exploring the demographic and social biases of existing unified MLLM models [321] is an important topic to ensure responsible deployment. As these models become increasingly capable across diverse modalities and tasks, unintentional amplification of cultural stereotypes, gender bias, or geographic imbalances embedded in pretraining data may result in harmful outputs. Future work should investigate effective fairness-aware training pipelines. Finally, enabling personalized knowledge-driven generation within unified MLLMs [322] is an emerging and important direction. Personalized models aim to incorporate user-provided concepts—such as specific objects, characters, or styles—into the model’s understanding and generation capabilities. However, current approaches often treat understanding and generation separately, using distinct concept embeddings for each task. This separation limits the model’s ability to generalize to compositional prompts that require implicit knowledge, such as generating ”$<$bo$>$ wearing its hat” without explicitly describing the hat. Unifying personalized understanding and generation under a shared modeling framework would allow better semantic grounding and contextual generalization.
To the best of our knowledge, most of current unified multimodal models primarily emphasize image understanding and text-to-image generation, while capabilities such as image editing are only attained through post-finetuning. Moreover, advanced functionalities like spatially controlled image generation, subject(s)-driven image generation, and interleaved image-text generation remain largely unexplored in the unified framework. Consequently, we believe there are abundant opportunities to advance the field by addressing key areas such as architectural design, training efficiency, dataset curation, evaluation methodologies, fairness, and reasoning to achieve unified multimodal models.